Use the following information about the labor market:

N=100-6W

N=28+3W

Y=A*K.5N.5

a)      Graph each of these curves. Make sure to completely label the graph and indicate which curve is the labor supply and which is labor demand.

b)      What is the market-clearing wage? Suppose that no worker is willing to work for less than $0. What are the employee and employer surpluses?

c)      Suppose that A=1 and K=4. Graph the production function. Given the number of workers you found in b) what is the output that is produced?

d)      Supposed that the government wants to help the poor. To do this they institute a minimum wage $1 higher than the wage you found in b). Show this graphically. How many workers will be hired? What are the employee and employer surpluses? Does this policy on average help employees? What will be the output produced?


1
Expert's answer
2021-01-15T18:20:36-0500

a)



b)


"100-6w=28+3w"


"w=8"


"N=52"


"\frac {1}{2}\times 52\times \frac {52}{6}=225"

c)



d)